OpenStudy (anonymous):

Workers Inc. just bought 10 boxes of pens to put in their storage room for employees to use as needed. Each box contains 30 pens. If each of the 28 employees working for Workers Inc. takes 4 pens when they are first brought into the office, how many pens will be left in the storage room? 272 296 160 188

OpenStudy (catlover5925):

i have answered this question before 28 employees take 4 pens so 28 * 4 = 112 now 10 boxes * 30 pens each box = 300 take 300 - 112= 188

OpenStudy (anonymous):

Lisa is on a trip of 2,753 miles. She has already traveled 218 miles. She has 5 days left in her trip. How many miles does she need to travel each day to complete the trip? 507 725 2,028 2,535 thanks

OpenStudy (catlover5925):

2,753 - 218 = 2535 now if she has 5 days left divide by 5 2535/5 = 507

OpenStudy (anonymous):

Robert delivers papers every morning. On Sundays he has to deliver 56 papers and on each of the other days he only has to deliver 31 papers. How many papers does Robert deliver each week? 242 273 211 392 thanks

OpenStudy (anonymous):

242

OpenStudy (anonymous):

For a scavenger hunt, Jim's mom distributed a bag of 306 jelly beans evenly into 17 plastic containers and hid them around the yard. If, after the hunt, Jim has a total of 144 jelly beans, then how many of the plastic containers did he find? 8 162 7 297

OpenStudy (catlover5925):

igtg sorry have @RussianBrain98 help